OpenStudy (anonymous):

Find the perimeter and area of a right triangle if the shortest side is 2.5 m. and the longest side is 6.5 m. Include correct units with each part of your solution.

OpenStudy (cwrw238):



OpenStudy (cwrw238):

find the length of the other side using pythagoras: x = sqrt( 6.5^2 - 2.5^2) area = 0.5 * 2.5 * x perimeter = 6.6 + 2.5 + x

OpenStudy (cwrw238):

follow that?

OpenStudy (anonymous):

 find x using b^2=c^2-b^2